# Quasi-Trefftz spaces for a first-order formulation of the Helmholtz equation

### v1

@@ Abstract
This work is the first step in the development of quasi-Trefftz methods for first-order differential systems. It focuses on discrete quasi-Trefftz spaces, starting from their definition and including construction of corresponding bases together with their computational aspect.
@@ 1 Introduction
In this sense, a given Trefftz space is more efficient than a polynomial space in approximating a smaller set of functions. Note that other methods, like Trefftz methods, make use of problem-dependent function spaces. These include Discontinuous-Petrov-Galerkin methods [BMS02, DG10], multiscale methods [AHP21, Pet17] and XFEM methods [AH08, SCB01, MB96].
@@
Quasi-Trefftz spaces have been investigated for scalar problems. To take advantage of the oscillating behavior, the first ansatz of quasi-Trefftz functions was wave-based, the idea being to add higher-degree polynomial terms in the phase of a plane wave. Other wave-based ansatz were also introduced, as well as purely a polynomial ansatz, see for instance [IGS25].
Beyond scalar problems, this article is concerned with the study of quasi-Trefftz spaces for a first-order system formulation of Helmholtz equation, as a stepping stone towards more general mixed-form problems.
@@ 1.3 Quasi-Trefftz properties and corresponding notation
In the case of a scalar equation of the form $\mathcal{L}u=0$, a given function $\varphi$ satisfies a quasi-Trefftz property for example if the Taylor expansion of its image by $\mathcal{L}$ is zero up to some order:
$\mathcal{T}_{q}\mathcal{L}\varphi=0.$
@@
    $\{\varphi:\bm{x}\mapsto\exp P(\bm{x}),P\in\mathbb{P}^{d},\mathcal{T}_{d-\gamma}\mathcal{L}\varphi=0\},$
    and this space has infinite dimension; these were first introduced in [IGD13] and further studied in [IG15];
@@
    $\{\varphi:\bm{x}\mapsto P(\bm{x})\exp(\mathrm{i}\omega\mathbf{k}\cdot\bm{x}),P\in\mathbb{P}^{d},\mathbf{k}\in\mathbb{C}^{2},|\mathbf{k}|=1,\mathcal{T}_{d-\gamma}\mathcal{L}\varphi=0\},$
    and this space also has infinite dimension; these were first introduced in [IG21] and further studied in [IGS25];
@@
    $\{\varphi:\bm{x}\mapsto P(\bm{x}),P\in\mathbb{P}^{d},\mathcal{T}_{d-\gamma}\mathcal{L}\varphi=0\},$
    and this is a finite-dimensional space; a recent comprehensive study of these [IG25].
@@ 2.1 Pressure-velocity definition
therefore the first two scalar quasi-Trefftz properties in the definition of ${\mathbb{S}}_{d}$ are in this case exact Trefftz properties. They can then compactly be written as $\mathrm{i}\omega\rho\bm{v}+\nabla p=0$.
@@ 2.2 Decoupled pressure
Given a point $\bm{x}_{0}\in\mathbb{R}^{2}$, a polynomial degree $d\in\mathbb{N}$ with $d\geq 2$, as well as the parameters of the partial differential operator $\mathcal{L}$, namely $\omega\in\mathbb{R}$, $\rho\in\mathbb{R}_{+}$, and $c\in\mathcal{C}^{d-2}$ at $\bm{x}_{0}$ with $c(\bm{x}_{0})\neq 0$, then the corresponding quasi-Trefftz space satisfies
@@ 2.3 Best approximation properties
Proving high-order convergence of Galerkin-type methods relies on best approximation properties of discrete spaces. In the case of quasi-Trefftz spaces, such properties can be proved via a well-chosen Taylor polynomial of the function to be approximated, as in the scalar case [IG25].
@@
Of course, in general, differential operators and Taylor truncation operators do not commute. Yet as we will see below, some such operators do commute when applied to smooth solution to the differential system.
@@
$\left\{\begin{array}[]{l}\|(p,v_{x},v_{y})(\bm{x})-(p_{a},v_{x,a},v_{y,a})(\bm{x})\|\leq C|\bm{x}-\bm{x}_{0}|^{d+1}\\ \|(\nabla p,\nabla v_{x},\nabla v_{y})(\bm{x})-(\nabla p_{a},\nabla v_{x,a},\nabla v_{y,a})(\bm{x})\|\leq C|\bm{x}-\bm{x}_{0}|^{d}\end{array}\right$.
for some constant $C$ independent of $\mathbf{x}$, but depending on $(p,v_{x},v_{y})$, $\mathbf{x}_{0}$ and $d$.
@@ 3.2 From the decoupled pressure
Here the contruction of the pressure and velocity can be completely decoupled. The first set of scalar differential equations has a solution at each iteration $k$, it falls under the theory developed in [IG25] so at each iteration $k$ the equation has a solution since the right-hand side would be known from previous iterations. Then the second set of equations provides an explicit solution. Altogether, this can be implemented as described in Algorithm 2.
@@ 3.3 Extension to 3D
As for Algorithm 2, it simply requires three-dimensional versions of the SolveLap and CompGrad functions. The latter is straightforward while the former can be found in [IG25].
@@ 6 Beyond Helmholtz and polynomial bases
In [IG25], a general framework developed to study polynomial quasi-Trefftz spaces for scalar PDEs was presented. In [FIG25], this general framework was extended to GPW quasi-Trefftz spaces for scalar PDEs. Similarly, it would be interesting to extend the present work beyond the polynomial ansatz to a GPW ansatz. In this case, the quasi-Trefftz space cannot be characterized as the kernel of a linear operator. However, it is expected that, as for the scalar case, it would be possible to split the non-linear system defining a quasi-Trefftz property for the system case into a hierarchy of linear subsystems according to the graded structure of polynomial spaces.

### v2

@@ Abstract
This work is concerned with the development of quasi-Trefftz methods for first-order differential systems. It focuses on discrete quasi-Trefftz spaces, starting from their definition and including the construction of corresponding bases together with their computational aspect.
This is the first attempt at constructing quasi-Trefftz bases for a problem governed by a first-order system without relying on an auxiliary scalar equation. A decoupling approach, with a second order scalar equation for the one unknown, is proposed here simply as a point of comparison to this new approach.
Acknowledgments
L.-M. Imbert-Gérard acknowledges support from the US National Science Foundation: this material is based upon work supported by the United States National Science Foundation under Grant No. DMS-2110407.
Statements and declarations
L.-M. Imbert-Gérard has disclosed an outside interest in Airbus Central R&T to the University of Arizona. Conflicts of interest resulting from this interest are being managed by The University of Arizona in accordance with its policies.
@@ 1 Introduction
In this sense, a given Trefftz space is more efficient than a polynomial space in approximating a smaller set of functions. Note that other methods, like Trefftz methods, make use of problem-dependent function spaces. These include Embedded Trefftz methods [CP23], Discontinuous-Petrov-Galerkin methods [BMS02, DG10], multiscale methods [AHP21, PET17] and XFEM methods [AH08, SCB01, MB96].
@@
Quasi-Trefftz spaces have been investigated for scalar problems. To take advantage of the oscillating behavior, the first ansatz of quasi-Trefftz functions was wave-based, the idea being to add higher-degree polynomial terms in the phase of a plane wave. Other wave-based ansatz were also introduced, as well as purely a polynomial ansatz, see for instance [IS25]. More recently in [IMB25a], polynomial quasi-Trefftz spaces and bases were tackled systematically for problems governed by scalar PDEs, thanks to an abstract linear algebra framework, for operators of arbitrary order ($\geq 1$) and in any dimension ($\geq 2$). This work relies on the fact that, in the scalar case, the principal part of the differential operator (i.e. terms of higher order derivatives with constant coefficients) is surjective between spaces of homogeneous polynomials with associated orders. By contrast, the system case is fundamentally different from the scalar case: for a vector-valued differential operator, the corresponding principal part is not necessarily surjective between appropriate spaces of homogeneous polynomials. Hence in this case the systematic approach developed for the scalar is not sufficient to handle even simple problems such as a first-order system for Helmholtz.
Both for time-domain and frequency-domain models, first-order formulations for wave propagation problems can be leveraged to develop new numerical methods efficient in the particularly challenging high-frequency regime [MP18, MC23, PES25, RPT24, RIV25]. Following this direction, this work is a stepping stone towards the development of quasi-Trefftz methods for more general problems governed by Friedrichs systems [FRI58, EG21]. It tackles a problem governed by a first-order formulation of the Helmholtz equation. While a quasi-Trefftz method for a first-order system formulation of the wave equation was already studied in [IMS23], there the notion of quasi-Trefftz space was only defined via a quasi-Trefftz space for an auxiliary second-order scalar equation. With a new approach, the present work is concerned with the study of quasi-Trefftz spaces for a first-order system formulation, but instead of relying on an auxiliary scalar equation, it focuses solely on the first-order system. This is crucial for further development of quasi-Trefftz methods, as a decoupling approach is not always possible for more general first-order systems.
@@ 1.3 Quasi-Trefftz properties and corresponding notation
In the case of a scalar equation of the form $\mathcal{D}u=0$, a given function $\varphi$ satisfies a quasi-Trefftz property for example if the Taylor expansion of its image by $\mathcal{D}$ is zero up to some order:
$\mathcal{T}_{q}\mathcal{D}\varphi=0.$
@@
    $\{\varphi:\bm{x}\mapsto\exp P(\bm{x}),P\in\mathbb{P}^{d},\mathcal{T}_{d-\gamma}\mathcal{D}\varphi=0\},$
    and this space has infinite dimension; these were first introduced in [ID13] and further studied in [IMB15];
@@
    $\{\varphi:\bm{x}\mapsto P(\bm{x})\exp(\mathrm{i}\omega\mathbf{k}\cdot\bm{x}),P\in\mathbb{P}^{d},\mathbf{k}\in\mathbb{C}^{2},|\mathbf{k}|=1,\mathcal{T}_{d-\gamma}\mathcal{D}\varphi=0\},$
    and this space also has infinite dimension; these were first introduced in [IMB21] and further studied in [IS25];
@@
    $\{\varphi:\bm{x}\mapsto P(\bm{x}),P\in\mathbb{P}^{d},\mathcal{T}_{d-\gamma}\mathcal{D}\varphi=0\},$
    and this is a finite-dimensional space; a recent comprehensive study of these [IMB25b].
@@ 2.1 Pressure-velocity definition
therefore the first two scalar quasi-Trefftz properties in the definition of ${\mathbb{S}}_{d}$ are in this case exact Trefftz properties. They can then be compactly written as $\mathrm{i}\omega\rho\bm{v}+\nabla p=0$.
@@ 2.2 Decoupled pressure
Given a point $\bm{x}_{0}\in\mathbb{R}^{2}$, a polynomial degree $d\in\mathbb{N}$ with $d\geq 2$, as well as the parameters of the Helmholtz partial differential operator $\mathcal{L}$ introduce in Section 1.1, namely $\omega\in\mathbb{R}$, $\rho\in\mathbb{R}_{+}$, and $c\in\mathcal{C}^{d-2}$ at $\bm{x}_{0}$ with $c(\bm{x}_{0})\neq 0$, then the corresponding quasi-Trefftz space satisfies
@@
In other words, even though the space ${\mathbb{S}}_{d}$ introduced in Definition 2.1.1 is defined directly from the quasi-Trefftz approximation of the first-order system, it can also be expressed from a quasi-Trefftz approximation of a decoupled Helmholtz equation for the pressure and a post-processing to get the velocity.
@@ 2.3 Best approximation properties
Proving high-order convergence of Galerkin-type methods relies on best approximation properties of discrete spaces. In the case of quasi-Trefftz spaces, such properties can be proved via a well-chosen Taylor polynomial of the function to be approximated, as in the scalar case [IMB25b].
@@
Of course, in general, differential operators and Taylor truncation operators do not commute, since for instance $\mathcal{T}_{0}\circ\partial_{x}\neq\partial_{x}\circ\mathcal{T}_{0}$ as the images of $f:\bm{x}=(x,y)\mapsto x$ are different ($1\neq 0$). Yet as we will see below, some such operators do commute when applied to smooth solution to the differential system. This relies on the fact that given $m\in\mathbb{N}$ and a multi-index $(\alpha_{1},\alpha_{2})\in\mathbb{N}_{0}^{2}$ with $\alpha_{1}+\alpha_{2}\leq m-1$ then $\mathcal{T}_{m-\alpha_{1}-\alpha_{2}}\circ[\partial_{x}^{\alpha_{1}}\partial_{y}^{\alpha_{2}}]=[\partial_{x}^{\alpha_{1}}\partial_{y}^{\alpha_{2}}]\circ\mathcal{T}_{m}$.
@@
$\left\{\begin{array}[]{l}\|(p,v_{x},v_{y})(\bm{x})-(p_{a},v_{x,a},v_{y,a})(\bm{x})\|_{\mathbb{C}^{3}}\leq C\|\bm{x}-\bm{x}_{0}\|_{\mathbb{C}^{3}}^{d+1}\\ \|(\nabla p,\nabla v_{x},\nabla v_{y})(\bm{x})-(\nabla p_{a},\nabla v_{x,a},\nabla v_{y,a})(\bm{x})\|_{\mathbb{C}^{3}}\leq C\|\bm{x}-\bm{x}_{0}\|_{\mathbb{C}^{3}}^{d}\end{array}\right$.
for some constant $C$ independent of $\mathbf{x}$, but depending on $(p,v_{x},v_{y})$, $\mathbf{x}_{0}$ and $d$, where $\|\cdot\|_{\mathbb{C}^{3}}$ is the euclidean norm on $\mathbb{C}^{3}$}.
@@ 3.2 From the decoupled pressure
Here the contruction of the pressure and velocity can be completely decoupled. The first set of scalar differential equations has a solution at each iteration $k$, it falls under the theory developed in [IMB25b] so at each iteration $k$ the equation has a solution since the right-hand side would be known from previous iterations. Then the second set of equations provides an explicit solution. Altogether, this can be implemented as described in Algorithm 2.
@@ 3.3 Extension to 3D
As for Algorithm 2, it simply requires three-dimensional versions of the SolveLap and CompGrad functions. The latter is straightforward while the former can be found in [IMB25b].
@@ 6 Beyond Helmholtz and polynomial bases
In [IMB25b], a general framework developed to study polynomial quasi-Trefftz spaces for scalar PDEs was presented. In [FI25], this general framework was extended to GPW quasi-Trefftz spaces for scalar PDEs. Similarly, it would be interesting to extend the present work beyond the polynomial ansatz to a GPW ansatz. In this case, the quasi-Trefftz space cannot be characterized as the kernel of a linear operator. However, it is expected that, as for the scalar case, it would be possible to split the non-linear system defining a quasi-Trefftz property for the system case into a hierarchy of linear subsystems according to the graded structure of polynomial spaces.